OpenStudy (anonymous):

The table below shows the values of f(n) for different values of n. n 1 2 3 4 5 6 f(n) 1 2 5 12 29 70 Which recursive function best represents the values shown in the table? (4 points) Question 12 options: 1) f(1) = 1, f(2) = 2, f(n) = 2f(n -1) f(n - 2); n > 2 2) f(1) = 1, f(2) = 2, f(n) = f(n -3) + f(n - 2); n > 2 3) f(1) = 1, f(2) = 2, f(n) = 2f(n -1) + f(n - 2); n > 2 4) f(1) = 1, f(2) = 2, f(n) = f(n -3) f(n - 2); n > 2

OpenStudy (anonymous):

can someone please help me wit this

OpenStudy (xapproachesinfinity):

the third option is correct since f(3)=5 , f(3)=2f(3-1)+f(3-2)=2f(2)+f(1)=2.2+1=5 check for f(4)=2f(4-1)+f(4-2)=2f(3)+f(2)=2.5+2=12 and so on

OpenStudy (anonymous):

Welcome to OpenStudy! :)